Can the sum of the magnitude of two vectors ever be equal to the magnitude of the sum of these two vectors???
 
 

Grade:12th pass

2 Answers

Anirudh
6 years ago
Mathematically,Let the vectors be a and b with magnitudes a and b respectively and the angle between them be x.Magnitude of the sum of a and b is√(a^2+b^2+2abcosxDifference in their magnitudes isa-bHence,√(a^2+ b^2+2ab cosx) = a-bSquaring both sides,a^2+b^2+2ab cos x = a^2+ b^2–2ab2ab cosx+2ab =02ab(cosx +1) =0Since 2ab can`t be zero,Cos x+1=0Cosx=-1X=180
Khimraj
6 years ago
Yes. this is posible but only when angle between these two vecor is zero.
Proof:-
Let first vector has magnitude A and second vector has magnitude B. And angle between these two is \Theta.
Then
A+B = (A2+B2+2ABcos\Theta)½
(A+B)2 = (A2+B2+2ABcos\Theta)
A2+B2+2AB = A2+B2+2ABcos\Theta
cos\Theta = 1
So \Theta = 0.
Hope it clears. If you like answer then please approve it.